OpenStudy (somegirl1999):

Find the area of the triangle. (1 point) 21.2 in2 10.5 in2 21 in2 42 in2

OpenStudy (somegirl1999):

i think the answer is 10.5 is tht right ?

OpenStudy (dls):

yes

OpenStudy (somegirl1999):

thank you :)

OpenStudy (anonymous):

The area of WHAT triangle? I see no triangle.

OpenStudy (somegirl1999):

Lol hold on

OpenStudy (anonymous):

exactly wat I was thinking

OpenStudy (somegirl1999):



OpenStudy (dls):

21

OpenStudy (anonymous):

yes 10.5 is the answer

OpenStudy (somegirl1999):

thank you so so much

OpenStudy (dls):

21 in^2

OpenStudy (anonymous):

\[A = \frac{ 1 }{ 2}bh\]

OpenStudy (dls):

\[\LARGE A=\frac{1}{2} \times base \times height\]

OpenStudy (dls):

\[\LARGE A=\frac{1}{2} \times 6 \times 7\]

OpenStudy (somegirl1999):

Thanx :P

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

lol